AB INITIO CALCULATION OF UO2

\underline{QIANG CHANG}, and RUSSELL M. PITZER, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Ab initio spin-orbit configuration interaction calculations are performed on the UO_{2} molecule,which has been observed in matrix isolation and laser ablation vacuum ultraviolet spectroscopic studies. We found the ground state to be 5f^{1}7s^{1} ^{3}Φ_{2u}. There are four excited states below 32,000 cm^{−1}that have large transition moments with the gound state. The highest of these excited state is 5f^{1}σ_{u}^{1} ^{3}Φ_{2g}. It has two minima in its potential curve. The σ_{u} orbital at the minimum with shorter bond distance has mostly 7p character, while it has mostly 5f character at the longer bond distance minimum. There are only a few dipole-allowed emission lines from this state to lower-energy states. Most electronic states in this region of the spectrum have g symmetry because most of the orbitals involved (5f,7s,7p) have the same (u) inversion symmetry.